OpenStudy (anonymous):

Determine whether the series converges absolutely, conditionally, or not at all Series of (sin(n) * pi)/sqrt(n) from n = 1 to infinity

OpenStudy (anonymous):

\[\sum_{1}^{\infty} \sin n * \pi / \sqrt{n}\]